OpenStudy (elenathehomeschooler):

In 1991, the cost of mailing a 1 oz. first-class letter was 29 cents, and the inflation rate was 4.6%. If the inflation rate stayed constant, the function C(t) = .29(1.046)t would represent the cost of mailing a first-class letter as a function of years since 1991. 2. In 2007, the cost of mailing a first-class letter was 41 cents. Has the inflation rate stayed constant since 1991? Explain.

OpenStudy (elenathehomeschooler):

@jim_thompson5910 can you help?

jimthompson5910 (jim_thompson5910):

2007 - 1991 = 16 so plug t = 16 into the C(t) function and tell me what you get

OpenStudy (elenathehomeschooler):

i got -15.40

jimthompson5910 (jim_thompson5910):

you should get some number between 0 and 1

OpenStudy (elenathehomeschooler):

0.03

jimthompson5910 (jim_thompson5910):

\[\Large C(t) = 0.29(1.046)^t\] \[\Large C(16) = 0.29(1.046)^{16}\] \[\Large C(16) = \underline{ \ \ \ \ \ \ \ \ \ }\] fill in the blank

jimthompson5910 (jim_thompson5910):

use a calculator to evaluate `0.29*(1.046)^16`

OpenStudy (elenathehomeschooler):

it will be 0.596

jimthompson5910 (jim_thompson5910):

so approx 60 cents if we use the C(t) function, the price should be 60 cents but it's actually 41 cents so the rate of inflation is NOT the same (it's NOT constant). The rate of inflation has gone down